OpenStudy (anonymous):

Help ASAP?? I just need answers checked for Algebra 2 I fan and medal

OpenStudy (anonymous):

1.What is the simplified form of (9x^2-25)/(3x+5) I answered 3x+5 with a restriction of x =/ 5/3 2. What is the simplified form of (12z^2-7z-12)/(3z^2+2z-8) I answered (4z+3)/(z+2) 3. What is the simplified form of {(x+3)/x^2-x-12}*{(x-4)/x^2-8x+6)} I answered 1/(x-4)^2

OpenStudy (cwrw238):

1 incorrect this is difference of 2 squares a^2 - b^2 = (a - b)( a + b) correct is 3x - 5 x not = 5/3

OpenStudy (cwrw238):

2 you neeed to factor top and bottom

OpenStudy (anonymous):

So would it be -5/3?

OpenStudy (cwrw238):

no x not = 5/3

OpenStudy (cwrw238):

- thats the restriction

OpenStudy (anonymous):

oh okay sorry i got confused

OpenStudy (cwrw238):

1. answer is 3x - 5 with restriction x not = 5/3

OpenStudy (anonymous):

okay, what about the other two?

OpenStudy (cwrw238):

12z^2-7z-12 - can you factor this?

OpenStudy (cwrw238):

= (4z + 3(3z - 4) what about factors of 3z^2+2z-8 ?

OpenStudy (cwrw238):

= (3z )(z ) can you fill in the blanks?

OpenStudy (cwrw238):

you need to factor -8

OpenStudy (anonymous):

Would it be (4z+3) (3z-4)

OpenStudy (cwrw238):

- thats the numerator - i did that in my above post what about the denominators factors?

OpenStudy (anonymous):

Oops sorry I didn't see that post. And the denominator factors would be (3z-4) and (z+2)?

OpenStudy (cwrw238):

right

OpenStudy (cwrw238):

3z - 4 cancels out so your answer to 2 is correct

OpenStudy (cwrw238):

for 3 you need to factor the denominators and see if any factor will cancel

OpenStudy (anonymous):

The factors for the first denominator are (x+3) and (x-4)

OpenStudy (cwrw238):

right - r u sure you've copied the other one correctly?

OpenStudy (anonymous):

Oops, I messed up the second denominator, it is supposed to be x^2-8x+16

OpenStudy (cwrw238):

oh ok what does that factor to?

OpenStudy (anonymous):

(x-4)^2

OpenStudy (cwrw238):

right

OpenStudy (cwrw238):

3 is correct

OpenStudy (anonymous):

Okay thank you. Could you help with a few more?

OpenStudy (cwrw238):

sorry gotta go

OpenStudy (anonymous):

Okay that's fine. Thank you!

OpenStudy (cwrw238):

yw